# Invariants for $\mathbb{G}_{(r)}$-modules

### v1

@@ 4. Exponential Jordan type
Assertion (1) follows from Proposition 4.5(1) and the formula of (4.4.1). Assertion (2) follows from Proposition 4.5(3). Assertions (3) and (4) and their proofs are direct analogues of Theorem 2.11 (3) and (4). ∎
@@
The reader might wish to compare $\Theta_{g,r}$ of (4.8.1), which is an element in the group algebra of $\mathbb{G}_{(r),k[\mathcal{C}_{r}(\mathfrak{g})]}$, with the image of $t^{p^{r}}$ as given in (4.4.1) which is a $k(\underline{B})[t]/t^{p^{r}}$-point of $\mathbb{G}_{(r)}$.
The following proposition is a mild generalization of a result of Sobaje [20, Prop 2.3] and is closely related to [10, Prop 4.2]. This proposition implies that $\Theta^{exp}_{\mathfrak{g},r}$ (much simpler than $\Theta_{\mathbb{G},r}$) determines the support theory $M\mapsto(\mathbb{P}V_{r}(\mathbb{G}))_{M}$ for finite dimensional $\mathbb{G}_{(r)}$-modules. See Corollary 4.13 below.
@@
Let $(\mathbb{G},\mathcal{E}_{r})$ be an affine group scheme of exponential type of height $r$, let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module of dimension $m$, and let $\underline{B}$ be a scheme-theoretic point of $\mathcal{C}_{r}(\mathfrak{g})$. Consider
$\Theta^{exp}_{\mathfrak{g},r}\ \equiv\ \sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s}\quad\in\quad k[\mathcal{C}_{r}(\mathfrak{g})]\otimes\mathbb{G}_{(r)},$
$\Theta^{exp}_{\mathfrak{g},r,\underline{B}}\ \equiv\ \sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}\ =\ \sum_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*}(u_{r-s-1})\quad\in\quad k(\underline{B})\otimes\mathbb{G}_{(r)}.$
For any scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathbb{G})$,
$JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \frac{m}{p}[p]\quad\iff\quad JT(\Theta_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \frac{m}{p}[p].$
Proof.
We repeat Sobaje’s proof of [20, Prop 2.3] to show that $\Theta_{\mathfrak{g},r,\underline{B}}-\Theta^{exp}_{\mathfrak{g},r,\underline{B}}$ is a linear combination of terms each of which is a product of two or more factors which are $p$-nilpotent and thus have $p-1$-st power 0. The proposition then follows from [6, Prop 8] (see also the more general [15, Thm 1.12]).
We may write $\Theta_{\mathfrak{g},r,\underline{B}}\ =\prod_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}$ as the composition
(×$(\times_{s=0}^{r-1}\ u_{r-s-1})\circ(\otimes_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*})\circ(1\otimes\Delta):\ k(\underline{B})\otimes k[\mathbb{G}_{(r)}]\ \to\ k(\underline{B})\otimes k[\mathbb{G}_{(r)}]^{\otimes r}$
$\quad\to\ k(\underline{B})\otimes k[\mathbb{G}_{a(r)}]^{\otimes r}\to k(\underline{B}).$
Sobaje shows that the result equals $\Theta^{exp}_{\mathfrak{g},r,\underline{B}}\ \equiv\ \sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}$ plus other terms which are scalar multiples of products of at least two factors of the form $\Theta^{exp}_{\mathfrak{g},r,|s,\underline{B}}$, factors which are $p$-nilpotent and pair-wise commuting. ∎
Remark 4.11.
By definition, $\Theta^{exp}_{\mathfrak{g},1}\ =\ \Theta_{\mathfrak{g},1}$. As observed in [20, Prop 2.3] (see also the proof of [23, Prop 6.5]), $\Theta^{exp}_{\mathfrak{g},2}\ =\ \Theta_{\mathfrak{g},2}$. On the other hand, for $r>2$, $\Theta^{exp}_{\mathfrak{g},r}\ \not=\ \Theta_{\mathfrak{g},r}$.
Remark 4.12.
The Jordan type $JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)$ is confusingly related to the Jordan type for a rational $\mathbb{G}$-module with $\mathbb{G}$ a linear algebraic group (as defined in [10, Defn:4.4]). For linear algebraic groups, the Jordan type is not defined in terms of the push-forward to $k\mathbb{G}$ along 1-parameter groups $\mathbb{G}_{a}\to\mathbb{G}$ of elements of $k\mathbb{G}_{a}$. A consequence of this is that the expression of $\Theta_{\mathfrak{g},r}^{exp}$ as $\sum_{s=0}^{r-1}(\mathcal{E}_{B_{s}})_{*}(u_{r-s-1})$ in Proposition 4.10 is replaced by $\sum_{s\geq 0}(\mathcal{E}_{B_{s}})_{*}(u_{s})$ for linear algebraic groups.
Proposition 4.10 has the following immediate corollary.
Corollary 4.13.
Let $(\mathbb{G},\mathcal{E}_{r})$ be an affine group scheme of exponential type of height $r$. For any $\mathbb{G}$-module $M$ of dimension $m$, the (affine) support variety $V_{r}(\mathbb{G})_{M}\ \subset V_{r}(G)$ is the image under the bicontinuous map $\Phi_{\mathfrak{g},r}$ of
$\{\langle\underline{B}\rangle\in\mathcal{C}_{r}(\mathfrak{g}):JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ \not=\ \frac{m}{p}[p]\}\quad\subset\quad\mathcal{C}_{r}(\mathfrak{g}).$
The proof of Proposition 4.10 (together with Theorem 1.12 of [15]) also proves the following corollary.
Corollary 4.14.
Assume the notation and hypotheses of Proposition 4.10. Assume that $\underline{B}\in\mathcal{C}_{r}(\mathfrak{g})$ is a generic point. Then for any finite dimensional $\mathbb{G}_{(r)}$-module,
$JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\quad=\quad JT(\Theta_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M).$
We next formulate the projectivization $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ of $\Theta_{\mathfrak{g},r}^{exp}$ and give the analogue of Theorem 2.14 for $\tilde{\Theta}^{exp}_{\mathfrak{g},r}$. We denote by $\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$ the weighted projective variety $\operatorname{Proj}\nolimits k[\mathcal{C}_{r}(\mathfrak{g})]$, parallel to the notation of $\mathbb{P}V_{r}(\mathbb{G})$ for $\operatorname{Proj}\nolimits k[V_{r}(\mathbb{G})]$. As in Theorem 2.14, we retain the notation and conventions introduced prior to Proposition 2.13 relating graded modules for a graded commutative ring $R$ and sheaves on $\operatorname{Proj}\nolimits R$.
Proposition 4.15.
@@
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. The operator $\Theta_{\mathfrak{g},r}^{exp}$ of Proposition 4.10 determines a global section
@@
$\tilde{\Theta}_{\mathfrak{g},r}^{exp}\quad\in\quad\Gamma(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
@@
of the coherent, locally free sheaf $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}$-modules which satisfies the following properties.
@@
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ determines the map of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}$-modules
    $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}:(k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M)^{\sim}\quad\to\quad(k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M)^{\sim}(p^{r-1}).$
@@
    For any non-zero divisior $h\in k[\mathcal{C}_{r}(\mathfrak{g})]_{p^{r-1}}$, the natural isomorphism of $(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})})_{U_{h}}$-modules
    $h^{-1}\otimes 1:(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})_{|U_{h}}\quad\stackrel{{\scriptstyle\sim}}{{\to}}\quad(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}\otimes k\mathbb{G}_{(r)})_{|U_{h}}$
@@
    $(\tilde{\Theta}_{\mathfrak{g},r}^{exp})_{|U_{h}}\ \in\ \Gamma(U_{h},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})\ =\ (k[\mathcal{C}_{r}(\mathbb{G})][1/h])_{p^{r-1}}\otimes k\mathbb{G}_{(r)}$
@@
    $(h^{-1}\otimes 1)\cdot\Theta_{\mathfrak{g},r}^{exp}\ \in\ k[\mathcal{C}_{r}(\mathbb{G})][1/h]\otimes k\mathbb{G}_{(r)}.$
@@
    If $\langle\underline{B}\rangle$ is a scheme-theoretic point of $\mathbb{P}\mathcal{C}_{r}(\mathbb{G})$ satisfying the condition that $h(\langle\underline{B}\rangle)\not=0$ and if $\underline{B}$ is a scheme theoretic point of $\mathcal{C}_{r}(\mathbb{G})$ projecting to $\langle\underline{B}\rangle$ with residue field $k(\langle\underline{B}\rangle)\subset k(\underline{B})$, then the restriction along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathbb{P}\mathcal{C}_{r}(\mathbb{G})$ of $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ is given by
    (4.15.1) $k(\underline{B})\otimes_{\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}}\tilde{\Theta}_{\mathfrak{g},r}^{exp}\quad=\quad c\cdot\sum_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*}(u_{r-s-1})\ \in\ k(\underline{B})\otimes k\mathbb{G}_{(r)}$
@@
    For any $\mathbb{G}_{(r)}$-module $M$, the restriction of $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}$ along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$ as in (3) is a non-zero scalar multiple of the image of
    $\sum_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*}(u_{r-s-1})$ in $End_{k(\underline{B})}(k(\underline{B})\otimes M)$.

### v2

@@ 4. Exponential Jordan type
Assertion (3) follows from Theorem 2.15(5). Assertions (4) and (5) and their proofs are direct analogues of Theorem 2.15(3),(4). ∎
As an immediate corollary of Theorem 4.8(1), we have the following.
Corollary 4.9.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. Consider a geometric point $\underline{B}$ of $\mathcal{C}_{r}(\mathfrak{g})$ with image $x=\Phi_{\mathfrak{g},r}(\underline{B})$. For any finite dimensional $\mathbb{G}_{(r)}$-module $M$,
$JT(\Theta_{\mathfrak{g},r,M})(\underline{B}))\quad=\quad JT(\Theta_{\mathbb{G},r,M})(x).$
In other words,
$JT(\Theta_{\mathfrak{g},r,M})(-)\ =\ JT(\Theta_{\mathbb{G},rM}(-)\circ\Phi_{\mathfrak{g},r}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\ \to\ \mathcal{Y},$
for any finite dimensional $\mathbb{G}_{(r)}$-module $M$.
We denote by $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ the weighted projective variety $\operatorname{Proj}\nolimits k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$, parallel to the notation of $\mathbb{P}V_{r}(\mathbb{G})$ for $\operatorname{Proj}\nolimits k[V_{r}(\mathbb{G})]$. We omit the proof of the following elementary lemma.
Lemma 4.10.
A geometric point of $\operatorname{Spec}\nolimits K\to\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ corresponds to an equivalence class $\langle\underline{B}\rangle$ of non-zero $r$-tuples $\underline{B}=(B_{0},\ldots,B_{r-1})$ of $p$-nilpotent, pair-wise commuting elements of $K\otimes\mathfrak{g}$, where $(B_{0},\ldots,B_{r-1})$ is equivalent to $(B_{0}^{\prime},\ldots,B_{r-1}^{\prime})$ if and only it there exists some $0\not=\alpha\in K$ such that $B_{i}^{\prime}=\alpha^{p^{i}}B_{i}$ for each $i\ 0\leq i<r$.
Under the bicontinuous homeomorphism $\Phi_{\mathfrak{g},r}$ of Lemma 4.7, the projection
$V_{r}(\mathbb{G}){\text{-}}\{0\}\to\mathbb{P}V_{r}(\mathfrak{g})$ corresponds to $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})){\text{-}}\{0\}\to\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ sending $\underline{B}$ to $\langle\underline{B}\rangle$.
We next formulate the projectivization $\tilde{\Theta}_{\mathfrak{g},r}$ of $\Theta_{\mathfrak{g},r}$ and give the analogue of Theorem 2.18 for $\tilde{\Theta}_{\mathfrak{g},r}$. As in Theorem 2.18, we retain the notation and conventions introduced prior to Proposition 2.17 relating graded modules for a graded commutative ring $R$ and sheaves on $\operatorname{Proj}\nolimits R$.
Proposition 4.11.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. The operator $\Theta_{\mathfrak{g},r}$ of Theorem 4.8 determines a global section
$\tilde{\Theta}_{\mathfrak{g},r}\quad\in\quad\Gamma(\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})),\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
of the coherent, locally free sheaf $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules which satisfies the following properties.
  1. (1)
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\tilde{\Theta}_{\mathfrak{g},r}$ determines the map of quasi-coherent $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules
    $\tilde{\Theta}_{\mathfrak{g},r,M}:(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)^{\sim}\quad\to\quad(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)^{\sim}(p^{r-1}).$
  2. (2)
    For any non-zero divisior $h\in k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]_{p^{r-1}}$, let $U_{h}$ denote the affine open subset $Spec(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][h^{-1}])_{0}\ \subset\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$. The natural isomorphism of $(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))})_{U_{h}}$-modules
    $h^{-1}\otimes 1:(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})_{|U_{h}}\quad\stackrel{{\scriptstyle\sim}}{{\to}}\quad(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}\otimes k\mathbb{G}_{(r)})_{|U_{h}}$
    sends
    $(\tilde{\Theta}_{\mathfrak{g},r})_{|U_{h}}\ \in\ \Gamma(U_{h},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})\ =\ (k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][1/h])_{p^{r-1}}\otimes k\mathbb{G}_{(r)}$
    to
    $(h^{-1}\otimes 1)\cdot\Theta_{\mathfrak{g},r}\ \in\ k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][1/h]\otimes k\mathbb{G}_{(r)}.$
  3. (3)
    If $\underline{B}$ is a scheme theoretic point of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ projecting to $\langle\underline{B}\rangle$ such that $h(\langle\underline{B}\rangle)\not=0$, then the restriction along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ of $\tilde{\Theta}_{\mathfrak{g},r}$ is given by
    (4.11.1) $k(\underline{B})\otimes_{\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}}\tilde{\Theta}_{\mathfrak{g},r}\quad=\quad c\cdot\prod_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1})\ \in\ k(\underline{B})\otimes k\mathbb{G}_{(r)}$
    for some $0\not=c\in k(\underline{B})$.
Proof.
By Proposition 2.17, $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ is a coherent, locally free sheaf of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules on $\mathbb{P}\mathcal{C}_{r}$. Proposition 4.5(3) implies that $(\Phi_{r|s})^{*}:k[V_{r}(\mathbb{G})]\to k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$ is a map of graded $k$-algebras. Thus, Theorem 2.15(5) and Theorem 4.8(1) tell us that $\Theta_{\mathfrak{g},r}\in k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes k\mathbb{G}_{(r)}$ is homogenous of degree $p^{r-1}$. Therefore, $\Theta_{\mathfrak{g},r}$ determines $\tilde{\Theta}_{\mathfrak{g},r}$ in $\Gamma(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})$.
The proofs of assertions (1), (2), and (3) are verified by evident analogues for $\tilde{\Theta}_{\mathfrak{g},r,}$ of the proofs given in the proof of Theorem 2.18 for the corresponding statements for $\tilde{\Theta}_{\mathbb{G},r,}$
We proceed to simplify $\Theta_{\mathfrak{g},r}$ by replacing the product in (4.11.1) by a sum.
Definition 4.12.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. We define
$\Theta^{exp}_{\mathfrak{g},r}\ \equiv\ \sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s}\quad\in\quad k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes k\mathbb{G}_{(r)}.$
Thus, for each $\underline{B}\in\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$
$\Theta^{exp}_{\mathfrak{g},r,\underline{B}}\ \equiv\ \sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}\ =\ \sum_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1})\quad\in\quad k(\underline{B})\otimes k\mathbb{G}_{(r)}.$
The justification of replacing $\Theta_{\mathfrak{g},r}$ by $\Theta_{\mathfrak{g},r}^{exp}$ is implicit in the following proposition, a reformulation of [12, Prop 4.3] (following a result of Sobaje [24, Prop 2.3]).
Proposition 4.13.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$, let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module, and let $\underline{a}=\sum_{i=1}^{p}a_{i}[i]$ be a Jordan type which is maximal among the set of Jordan types $\{JT(\Theta_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M);\ \underline{B}\in\mathcal{C}_{r}(\mathbb{G})\}$ for $M$. For any scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$,
$JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \underline{a}\quad{\text{i}f\ and\ only\ if}\quad JT(\Theta_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \underline{a}.$
Proof.
In [24, Prop 2.3], Sobaje regards $\Theta_{\mathfrak{g},r,\underline{B}}\ =\prod_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}$ as the composition
$(\prod_{s=0}^{r-1}\circ(\otimes_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*})\circ(1\otimes\Delta):\ k(\underline{B})\otimes k\mathbb{G}_{a(r)}\ \to\ k(\underline{B})\otimes k\mathbb{G}_{a(r)}^{\otimes r}$
$\quad\to\ k(\underline{B})\otimes k\mathbb{G}_{(r)}^{\otimes r}\to k(\underline{B})\otimes k\mathbb{G}_{(r)}$
applied to $1\otimes u_{r-1}\in k(\underline{B})\otimes k\mathbb{G}_{a(r)}$. Using this composition, he verifies that $\Theta_{\mathfrak{g},r,\underline{B}}$ equals $\sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}$ plus other terms which are scalar multiples of products of at least two factors of this sum, factors which are $p$-nilpotent and pair-wise commuting. The proposition then follows from [17, Thm 1.12] (see also [8, Prop 8] and [27, Lem 6.4]). ∎
As shown in Theorem 2.23 and remarked in Remark 2.26, the support variety and its generalizations for a finite dimensional $\mathbb{G}_{(r)}$-module $M$ are determined by $JT_{\mathbb{G},r,M}(-)$ and thus by $JT_{\mathfrak{g},r,M}(-)$. (See Corollary 4.9.)
Since these support varieties are formulated in terms of maximality of Jordan types, Proposition 4.13 has the following corollary.
Corollary 4.14.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$ and let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module. Then the support variety and generalized support varieties of $M$ interpreted as closed subvarieties of $\mathcal{C}_{r}(\mathfrak{g})$ using $JT_{\mathbb{G},r,M}(-)$ are left unchanged by replacing $JT_{\mathbb{G},r,M}(-)$ by $JT^{exp}_{\mathbb{G},r,M}(-)$.
@@
By definition, $\Theta^{exp}_{\mathfrak{g},1}\ =\ \Theta_{\mathfrak{g},1}$. As observed in [24, Prop 2.3] (see also the proof of [27, Prop 6.5]), $\Theta^{exp}_{\mathfrak{g},2}\ =\ \Theta_{\mathfrak{g},2}$. On the other hand, for $r>2$, $\Theta^{exp}_{\mathfrak{g},r}\ \not=\ \Theta_{\mathfrak{g},r}$.
We next give the evident definition of the “exponential Jordan type function” of a $\mathbb{G}_{(r)}$-module, parallel to the definition of $JT_{\mathbb{G},r,M}(-)$ given in Definition 2.16.
Definition 4.16.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$ and let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module. We define
$JT^{exp}_{\mathfrak{g},r,M}(-):\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\quad\to\quad\mathcal{Y}$
by sending a geometric point $\underline{B}:\operatorname{Spec}\nolimits k(\underline{B})\to\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ to
(4.16.1) $JT(\Theta^{exp}_{\mathfrak{g},r,M})(\underline{B})\quad=\quad JT(\sum_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1}),k(\underline{B})\otimes M))\ ,$
the Jordan type of the operator $\Theta^{exp}_{\mathfrak{g},r,\underline{B}}$ on $k(\underline{B})\otimes M$.
Observe that the operator $(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1})$ on $k(\underline{B})\otimes M$ can be identified with the operator $1\otimes u_{r-s-1}\in k(\underline{B})\otimes k\mathbb{G}_{a(r)}$ on $(\mathcal{E}_{r|s,B_{s}})^{*}(k(\underline{B})\otimes M)$.
We next formulate the projectivization $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ of $\Theta_{\mathfrak{g},r}^{exp}$ and give the analogues of Theorem 2.18 for $\tilde{\Theta}_{\mathbb{G},r}$ and Proposition 4.11 for $\tilde{\Theta}_{\mathfrak{g},r}$.
@@
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. The operator $\Theta_{\mathfrak{g},r}^{exp}$ of Definition 4.12 determines a global section
@@
$\tilde{\Theta}_{\mathfrak{g},r}^{exp}\quad\in\quad\Gamma(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
@@
of the coherent, locally free sheaf $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules which satisfies the following properties.
@@
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ determines the map of quasi-coherent $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules
    $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}:(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)^{\sim}\quad\to\quad(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)^{\sim}(p^{r-1}).$
@@
    For any non-zero divisior $h\in k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]_{p^{r-1}}$, let $U_{h}$ denote the affine open subset $Spec(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][h^{-1}])_{0}\ \subset\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$. The natural isomorphism of $(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))})_{U_{h}}$-modules
    $h^{-1}\otimes 1:(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})_{|U_{h}}\quad\stackrel{{\scriptstyle\sim}}{{\to}}\quad(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}\otimes k\mathbb{G}_{(r)})_{|U_{h}}$
@@
    $(\tilde{\Theta}_{\mathfrak{g},r}^{exp})_{|U_{h}}\ \in\ \Gamma(U_{h},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})\ =\ (k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][1/h])_{p^{r-1}}\otimes k\mathbb{G}_{(r)}$
@@
    $(h^{-1}\otimes 1)\cdot\Theta_{\mathfrak{g},r}^{exp}\ \in\ k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][1/h]\otimes k\mathbb{G}_{(r)}.$
@@
    If $\underline{B}$ is a scheme theoretic point of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ projecting to $\langle\underline{B}\rangle$ such that $h(\langle\underline{B}\rangle)\not=0$, then the restriction along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ of $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ is given by
    (4.17.1) $k(\underline{B})\otimes_{\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}}\tilde{\Theta}_{\mathfrak{g},r}^{exp}\quad=\quad c\cdot\sum_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*}(1\otimes u_{r-s-1})\ \in\ k(\underline{B})\otimes k\mathbb{G}_{(r)}$
@@
    For any $\mathbb{G}_{(r)}$-module $M$, the restriction of $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}$ along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ as in (3) is a non-zero scalar multiple of the image of
    $\sum_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1})$ in $End_{k(\underline{B})}(k(\underline{B})\otimes M)$.